# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ Abstract
Machine learning has achieved remarkable success in many applications. However, existing studies are largely based on the closed-world assumption, which assumes that the environment is stationary, and the model is fixed once deployed. In many real-world applications, this fundamental and rather naive assumption may not hold because an open environment is complex, dynamic, and full of unknowns. In such cases, rejecting unknowns, discovering novelties, and then incrementally learning them, could enable models to be safe and evolve continually as biological systems do. This paper provides a holistic view of open-world machine learning by investigating unknown rejection, novel class discovery, and class-incremental learning in a unified paradigm. The challenges, principles, and limitations of current methodologies are discussed in detail. Finally, we discuss several potential directions for future research. This paper aims to provide a comprehensive introduction to the emerging open-world machine learning paradigm, to help researchers build more powerful AI systems in their respective fields, and to promote the development of artificial general intelligence.
@@ 1 Introduction
Artificial intelligence coupled with machine learning techniques is broadly used in many fields such as medical treatment [1], industry [2], transportation and scientific discovery [3]. Typically, supervised machine learning involves isolated classification or regression task, which learns a function (model) $f:\mathcal{X}\rightarrow\mathcal{Y}$ from a training dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}^{N}_{i=1}$ containing pairs of feature vector and ground-truth label [4]. Then, the model $f$ can be deployed to predict future encountered inputs. However, the current success of machine learning is largely based on the closed-world assumption [5, 6, 7], where the important factors of learning are limited to what has been observed during training. In the classification task, all the classes $y$ that the model will encounter during deployment must have been seen in training, i.e., $y\in\mathcal{Y}$. This assumption is often reasonable in restricted scenarios where possible classes are well-defined and unlikely to change over time. For example, in a handwritten digit recognition task, the closed-world assumption holds because the set of digits (0-9) is fixed and known in advance. Besides, this assumption also makes the data collection process easier and straightforward.
However, real-world applications often involve dynamic and open environments, where unexpected situations inevitably arise and instances belonging to unknown classes ($y\notin\mathcal{Y}$) may appear [8, 9]. For example, in non-stationary environments, a self-driving car may encounter novel objects that have been never learned before; myriad novel categories would emerge continually in web usage and face recognition systems. The closed-world assumption can be problematic in such situations. Firstly, models are overconfident and predict unknowns as training classes without hesitation [10, 11, 12], which can cause various harms from financial loss to injury and death. Secondly, models fail to extrapolate to novel classes by discovering and clustering them [13]. Thirdly, learning new stream data leads to catastrophic forgetting of previous knowledge [14]. To learn in such an endless variety of ever-changing scenarios, we need open-world learning to overcome these limitations by accommodating the dynamic and uncertain nature of real-world data. In this paradigm, a model is equipped to identify and reject inputs that deviate from training classes to keep safe, and then discovers new classes from unknowns and incrementally learns them to accumulate knowledge without re-training the whole model from scratch.
The general life cycle of an open-world learning (OWL) paradigm is illustrated in Figure 1. This process mainly consists of three key steps. The first step is unknown rejection, which requires the model to recognize test instances that belong to seen classes while also being able to detect or reject misclassified and unknown instances that do not belong to the training classes based on reliable confidence estimation [11, 15]. The second step is novel class discovery [13], which clusters the collected unknown samples in the buffer automatically based on the knowledge learned in the past. Finally, when the discovered classes have sufficient data, the system must extend the original multi-class classifier to incorporate new classes without retraining from scratch or catastrophic forgetting of previously learned knowledge [16, 17, 18]. By integrating unknown rejection, novel class discovery, and continual learning, the system is able to scale and adapt to an ever-evolving environment. In other words, the model can be aware of what it does not know and learns interactively after deployment (on the job) in the open world like humans.
In this article, we present a systematic review of recent research advances in open-world machine learning, with an emphasis on techniques concerning unknown rejection, novel class discovery, and class-incremental learning. The principles and limitations of current methods and the relation among them are discussed in detail. Finally, the possible challenges, research gaps, and outlook for the future development of open-world machine learning are presented. Our broad and insightful review will be helpful for researchers to apply this new learning paradigm in their own domains, and also be a call to build human-like, truly intelligent systems.
@@
Figure 1: Illustrations of the life cycle of a learning system in the open-world applications. (A) Humans continually learn new knowledge throughout their lives and maintain/use the previous knowledge, becoming increasingly smarter and more skillful over time. (B) Open-world machine learning aims to build a human-like system that can transfer and consolidate knowledge incrementally during deployment. (C) An open-world learning paradigm mainly includes three parts, i.e., unknown rejection, novel class discovery (NCD) and class-incremental learning (CIL).
2 Background
2.1 Problem Formulation and Notations
@@
Here, we give a holistic problem formulation and notations of open-world machine learning, comprising three tasks: unknown rejection, novel class discovery and class-incremental learning. Let $\mathcal{D}_{\text{train}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}_{\text{train}}$ denote the training dataset, and $\mathcal{D}_{\text{test}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}\subset\mathcal{X}\times\mathcal{Y}_{\text{test}}$ denote the test dataset. By default, the training data only contains labeled data from old and known classes, i.e., $\mathcal{C}_{\text{old}}=\mathcal{Y}_{\text{train}}$. We represent a deep neural network (DNN) based model $f$ with two components: a feature extractor $h$ and a unified classifier $g$, i.e., $f=g\circ h$. In the canonical closed-world setting, the test dataset shares categories with training data with no additional classes, i.e., $\mathcal{Y}_{\text{test}}=\mathcal{Y}_{\text{train}}$. By contrast, in the open-world, samples from new classes emerge, i.e., $\mathcal{Y}_{\text{train}}\neq\mathcal{Y}_{\text{test}}$, we further express it as $\mathcal{Y}_{\text{train}}\subset\mathcal{Y}_{\text{test}}$. In this paper, we not only consider how to classify the seen and known classes, but also focus on how to handle samples from new classes $\mathcal{C}_{\text{new}}=\mathcal{Y}_{\text{test}}\setminus\mathcal{Y}_{\text{train}}$, which is divided into three progressive ways: unknown rejection, novel class discovery and class-incremental learning. In unknown rejection, models only need to reject unknowns, while in novel class discovery and class-incremental, models are further required to classify/cluster samples from new classes. Considering the clustering objective, in novel class discovery, $\mathcal{D}_{\text{train}}$ could also contain some unlabeled data from $\mathcal{C}_{\text{new}}$. While in class-incremental learning, the training scheme could be divided into several stages, models are trained on streaming data, and we add a superscript $t$ to the dataset $\mathcal{D}_{\text{train}}^{t}$ and $\mathcal{D}_{\text{test}}^{t}$ to indicate each incremental step. Considering the differences among the three tasks, we will also give task-specific notations for three tasks when discussing them.
2.2 Canonical Use Cases
@@
In this section, we discuss several canonical applications of open-world machine learning. An illustration is provided in Figure 2.
@@
Autonomous Driving. Self-driving cars in real-world scenarios can encounter a nearly infinite number of unexpected and novel objects [19], and open-world learning can provide them with the adaptability needed to navigate safely and effectively. Specifically, when encountering a novel object on the road, e.g. a deer or even a hole on the road that does not appear in the training class set, the car should stop or slow down. This involves leveraging an unknown rejection algorithm that provides reliable confidence scores. If the confidence score is lower than a threshold, the system should raise an alarm so that the car takes a predefined safety action. Then, novel class discovery techniques are used to create the class identity for the unknown, new object. Finally, based on incremental learning methods, the car can continue to learn from a single or few samples of the object, enabling autonomous systems to quickly adapt to unexpected new environments without forgetting the skills already mastered. In conclusion, open-world learning allows the vehicle to learn from dynamic real-world environments and adapt its decision-making process, improving its driving behavior over time.
@@
Medical Diagnosis. Real-time medical applications require the AI system to address many heterogeneous problems that involve multiple, complex tasks. For example, a pulmonary model trained on existing lung data would not be able to reliably predict newly emerged pulmonary diseases like COVID-19, and the medical diagnostic tools can easily make mistakes [20]. In such cases, the model should be equipped with unknown rejection ability to output reliable and human-interpretable confidence scores along with its predictions, which helps doctors or participants to make safe decisions. For example, the input should be handed over to human doctors when the confidence of a disease diagnosis is low. When encountering more and more unknowns, the medical diagnosis system is supposed to continually learn from those unknowns or more data to improve its performance progressively [21, 22]. Since manual image grading is time-consuming, novel class discovery methods would largely accelerate the processing of labeling new data. The above adaptive capability ensures that medical diagnosis systems can learn from on-the-job experiences and apply this knowledge to similar situations in the future, which is very similar to the ways that human clinicians learn. In conclusion, open-world learning enables medical diagnoses to make more reliable clinical decisions and be more responsive to evolving environments.
Figure 2: Example applications of open-world machine learning systems such as (A) autonomous driving, (B) medical diagnosis, and (C) AI chatbot.
@@
AI Chatbot. With the launch of large language models like ChatGPT [23], AI chatbots have become widely used in many scenarios to help with daily tasks for users. In practice, it is critical for a chatbot to know what it doesn’t know. Specifically, if the chatbot does not know the answer, it is better to reject to response rather than provide incorrect information that may cause very serious mistakes in fields like clinical and legal matters [24, 25]. One way to avoid providing misleading information is to abstain from making a prediction based on unknown rejection methods. Further, if the user provides the correct answer or more data is available, the chatbot should improve its conversational skills continually without forgetting previously learned knowledge. A typical example is the hotel guest-greeting bot [5], where a chatbot greets known hotel guests, detects new guests, and then incrementally learns them by asking their names and taking photos. As can be seen, open-world learning makes the AI Chatbot more knowledgeable and powerful over time.
2.3 Overall Challenges of Open-world Learning

### v3

# Open-world Machine Learning: A Systematic Review and Future Directions

@@ Abstract
Machine learning has achieved remarkable success in many applications. However, existing studies are largely based on the closed-world assumption, which assumes that the environment is stationary, and the model is fixed once deployed. In many real-world applications, this fundamental and rather naive assumption may not hold because an open environment is complex, dynamic, and full of unknowns. In such cases, rejecting unknowns, discovering novelties, and then continually learning them, could enable models to be safe and evolve continually as biological systems do. This article presents a holistic view of open-world machine learning by investigating unknown rejection, novelty discovery, and continual learning in a unified paradigm. The challenges, principles, and limitations of current methodologies are discussed in detail. Furthermore, widely used benchmarks, metrics, and performances are summarized. Finally, we discuss several potential directions for further progress in the field. By providing a comprehensive introduction to the emerging open-world machine learning paradigm, this article aims to help researchers build more powerful AI systems in their respective fields, and to promote the development of artificial general intelligence.
Index Terms:
Machine learning, deep learning, open-world, unknown rejection, novel class discovery, continual learning.
@@ I Introduction
Artificial intelligence, coupled with machine learning techniques [1, 2], is broadly used in many fields, such as medical treatment [3], industry [4], transportation and scientific discovery [5]. Typically, supervised machine learning involves isolated classification or regression task, which learns a function (model) $f:\mathcal{X}\rightarrow\mathcal{Y}$ from a training dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}^{N}_{i=1}$ containing pairs of feature vector and ground-truth label [6]. Then, the model $f$ can be deployed to predict future encountered inputs. However, the current success of machine learning is largely based on the closed-world assumption [7, 8, 9], where the important factors of learning are limited to what has been observed during training. In the classification task, all the classes $y$ that the model will encounter during deployment must have been seen in training, i.e., $y\in\mathcal{Y}$. This assumption is often reasonable in restricted scenarios where possible classes are well-defined and unlikely to change over time. For example, in a handwritten digit recognition task, the closed-world assumption holds because the set of digits (0-9) is fixed and known in advance. Besides, this assumption also makes the data collection process easier and straightforward.
Figure 1: Illustration of open-world machine learning scenarios. In intelligent driving, robot manipulation, medical diagnosis and Chatbot scenarios, the environments are open, dynamic and complex. Open-world machine learning enables the model to learn and evolve from streaming data safely.
@@
Figure 2: Illustrations of the life cycle of a learning system in the open-world applications. (A) Humans continually learn new knowledge throughout their lives and maintain/use previous knowledge, becoming increasingly smarter and more skillful over time. (B) Open-world machine learning aims to build a human-like system that can transfer and consolidate knowledge continually during deployment. (C) An open-world learning paradigm mainly includes three parts, i.e., unknown rejection, novel class discovery and continual learning.
TABLE I: A comprehensive taxonomy of open-world machine learning, including all three tasks of unknown rejection, novel class discovery and continual learning. Each task is divided into several subcategories with representative methods.
Categories Representative Methods
Unknown Rejection Out-of-distribution Detection Post-hoc Inference Methods MSP [10], ODIN [11], Energy Score [12], Mahalanobis [13], ViM [14], WDiscOOD [15], KNN [16], GEN [17], Cosine [18], ASH [19], SCALE [20], MOS [21], Maxlogits [22]
Training-stage Methods SSL [23], LogitNorm [24], FMFP [25], classAug [26], AoP [27], OVA-PL[28], RCL [29], GenData [30], PixMix [31]
Outlier-aided Methods OE [32], OpenMix [33], VOS [34], POE [35], LoCoOp [36], NegPrompt [37], NegLabel [38], LocalPrompt [39], NPOS [40]
Open-Set Recognition Discriminative Models OpenMax [41], PROSER [42], C2AE [43], RAD [44], ODL [45]
Generative Models G-OpenMax [46], OpenGAN [47]
Hybrid Models OpenHybrid [48], CPN [49], CAC [50], OVA-PL[28], ARPL [51], CSSR [52]
Novel Class Discovery Novel Category Discovery (NCD) Isolated-training Methods KCL [53], MCL [54], DTC [55], ResTune [56]
Joint-training Methods RandStats [57, 58], NCL [59], OpenMix [60], UNO [61], DualRS [62], ComEx [63], IIC [64], rKD [65], TIDA [66], SCKD [67]
Continual NCD FRoST [68], NCDwF [69], KTRFR [70], ADM [71]
Generalized Category Discovery (GCD) Non-parametric Classification GCD [72], XCon [73], DCCL [74], GPC [75], PromptCAL [76], InfoSieve [77], CiPR [78], CMS [79]
Parametric Classification ORCA [80], PIM [81], SimGCD [82], $\mu$GCD [83], SPTNet [84], LegoGCD [85], ProtoGCD [86]
Continual GCD GM [87], IGCD [88], PA-GCD [89], MetaGCD [90], PromptCCD [91], Happy [92]
Extended Settings Other Learning Paradigms Active Learning (ActiveGCD [93]), Federated Learning (FedGCD [94]) Domain Adaptation (HiLo [95])
On-the-fly Discovery SMILE [96], PHE [97]
Fine-grained Discovery RAPL [98], SelEx [99], PHE [97], FineR [100]
Long-tailed Discovery ImbaGCD [101], Long-tailed NCD [102], Long-tailed GCD [103], BYOP [104], DA-GCD [105]
Multimodal Discovery SCD [106], CLIP-GCD [107], TextGCD [108], GET [109], FineR [100], CPT [110]
Continual Learning Regularization Method Explicit Weight Constraint EWC [111], SI [112], MAS [113], RWalk [114], VCL [115], UCL [116], KCL [117], AGS-CL [118], OWM [119], OGD [120], AOP [121], GPM [122], Adam-NSCL [123], NCL [124], CAF [125]
Implicit Knowledge Distillation LwF [126], EBLL [127], LwM [128], DMC [129], M2KD [130], MUC [131], CCIL [132], UCIR [133], PODNet [134], DER [135], DER++ [135], GeoDL [136], DDE [137], FOSTER [138]
Data Replay Methods Imbalance Calibration EEIL [139], UCIR [133], GDumb [140], iCaRL [141], BiC [142], WA [143], ScaIL [144], IL2M [145], SS-IL [146], ItO [147]
Generative Data Replay DGR [148], MeRGAN [149], FearNet [150], DMC [129], ABD [151], ILCAN [152], DGM [153], DiffClass [154], DDGR [155], SDDGR [156]
Feature Replay Methods Real / Generative Feature REMIND [157], LR [158], GFR [153], BI-R [159], FAdapt [160]
Prototype Replay SDC [161], PASS [162], PASS++ [163], SSRE [164], IL2A [165], Fusion [166], FeTrIL [167], FeCAM [168], SOPE [169], NAPA-VQ [170], ADC [171], LDC [172], DPCR [173], EFC [174], EFC++ [175], AdaGauss [176], RanPAC [177], SDCLoRA [178], PRL [179], Semi-IPC [180], FCS [181]
Model Expansion Methods Backbone Expansion P&C [182], AANets [183], DER [184], Piggyback [185], HAT [186], PackNet [187], DyTox [188]
Parameter Efficient Expansion L2P [189], CODA-Prompt [190], DualPrompt [191], SDCLoRA [178], PILoRA [192], SD-LoRA [193], Online-LoRA [194], LAE [195], InfLoRA [196], HiDe-Prompt [197], C-LoRA [198],C-CLIP [199], TSVD [200], CAPrompt [201], DESIRE [202], FM-LoRA [203], LGCL [204], PGP [205], EvoPrompt [206], RCS-Prompt [207], MISA [208], LORA-DRS [209], EASE [210], NoRGa [211], SAFE [212], VQ-Prompt [213], VPT-NSP [214], RAPF [215], C-ADA [216], PromptFusion [217]
Extended Settings Unsupervised Continual Learning FRoST [68], NCDwF [69], GM [87], IGCD [88], PA-GCD [89], PromptCCD [91], Happy [92]
Multimodel Large Language Model O-LoRA [218], HiDe-LLaVA [219], ModalPrompt [220], MOELoRA [221], Continual-LLaVA [222], DISCO [223], LLaCA [224], LLaVA-CMoE [225], LOIRE [226], FV [227], Adapt-$\infty$ [228], SAPT [229], ProgPrompts [230], DAS [231]
However, as illustrated in Fig. 1, real-world applications often involve dynamic and open environments, where unexpected situations inevitably arise and instances belonging to unknown classes ($y\notin\mathcal{Y}$) may appear [232, 233]. For example, in non-stationary environments, a self-driving car may encounter novel objects that have never been learned before; myriad novel categories would emerge continually in web usage and face recognition systems. The closed-world assumption can be problematic in such situations. Firstly, models are overconfident and predict unknowns as training classes without hesitation [10, 26, 234], which can cause various harms from financial loss to injury and death. Secondly, models fail to extrapolate to novel classes by discovering and clustering them [58] from various unknowns. Thirdly, learning new stream data leads to catastrophic forgetting of previous knowledge [235]. To learn in such an endless variety of ever-changing scenarios, we need open-world learning to overcome these limitations by accommodating the dynamic and uncertain nature of real-world data. In this paradigm, a model is equipped to identify and reject inputs that deviate from training classes to keep safe, and then discovers meaningful new classes from unknowns and continually learns them to accumulate knowledge without re-training the whole model from scratch.
The general life cycle of an open-world learning (OWL) paradigm is illustrated in Fig. 2, mainly consisting of three key steps. The first step is unknown rejection, which requires the model to recognize test instances that belong to seen classes while also being able to detect or reject misclassified and unknown instances that do not belong to the training classes based on reliable confidence estimation [26, 25]. The second step is novel class discovery [58], which clusters the collected unknown samples and identifies meaningful categories in the buffer automatically based on the knowledge learned in the past. Finally, when the discovered classes have sufficient data, the system must extend the original multi-class classifier to incorporate new classes with continual learning techniques, avoiding retraining from scratch or catastrophic forgetting of previously learned knowledge [236, 237, 238]. By integrating unknown rejection, novel class discovery, and continual learning, the system is able to scale and adapt to an ever-evolving environment. In other words, the model can be aware of what it does not know and learns interactively after deployment (on the job) in the open world like humans.
OWL has garnered significant attention and exploration in recent years, resulting in the emergence of numerous methodologies in this research field. The exploration of OWL traces its conceptual roots to early investigations into open-set recognition and continual learning, with foundational works around the 2010s emphasizing statistical and shallow models to detect unknowns or adapt to incremental data [232, 235]. An example is the introduction of open-set recognition frameworks [239], which aimed to reject unseen categories by quantifying uncertainty in decision boundaries, though these methods often struggled with high-dimensional real-world data. The advent of deep learning [240, 241] marked a paradigm shift, as neural networks demonstrated superior capability in representation learning. Notably, Hendrycks et al. [10] established a baseline using deep networks to detect out-of-distribution samples, highlighting the role of softmax confidence scores in identifying unknowns. This era also witnessed growing attention to catastrophic forgetting [235], spurring research into continual learning techniques to preserve historical knowledge while integrating new information. In recent years, pre-training models have been leveraged to enhance the reliability and adaptation of AI systems in open environments [25, 233]. These advancements underscore a shift toward end-to-end systems capable of self-awareness and autonomous knowledge expansion. Looking ahead, OWL remains a vibrant yet challenging frontier. Future research is to explore OWL with foundation models, a broader spectrum of data modalities, and neuromorphic mechanisms to achieve human-like adaptability in ever-evolving environments.
Although prior surveys [232, 242, 234] have reviewed some aspects of OWL, significant gaps persist in terms of timeliness, taxonomy and comprehensiveness, i.e., they are either relatively early reviews lacking discussions on the up-to-date advances, or limited to only a few aspects of OWL. Specifically, Zhou [232] briefly introduced the concept of emerging new classes, decremental and incremental features, as well as changing data distributions in OWL. Parmar et al. [242] mainly focused on open-set classification, and Zhang et al. [234] offered a discussion of scenarios where the model needs to learn to reject. However, these works focus mainly on early research and a specific aspect of OWL, while important methods developed in the last five years, which have significantly shaped the field, have yet to be included. To overcome the aforementioned limitations and fill the gap in the literature of OWL, we present this comprehensive survey with a broader spectrum of learning tasks of OWL, including all three steps in Fig. 2 (c). For each task, we further provide a more detailed sub-taxonomy with up-to-date research advances. We undertake a comprehensive review of over four hundred articles, introducing a taxonomy delineating unknown rejection, novel class discovery, and continual learning. As shown in Table I, for each sub-topic, we divide existing methods into overarching categories and further into several subcategories. The principles and limitations of current methods and the relationship among them are discussed. Besides, existing theoretical results are also presented. We also summarize the commonly used datasets and evaluation metrics, and systematically evaluate different methods across various benchmarks. Furthermore, the possible challenges, research gaps, and potential research directions for the future development of open-world machine learning are presented. This survey updates the field’s narrative, being helpful for researchers to apply this new learning paradigm in their own domains, and also calls for building human-like, truly intelligent systems.
The rest of this paper is organized as follows: Section II provides the basic problem formulation, notations, several canonical use cases, and an overview of open-world machine learning. Section III presents unknown rejection, which requires the model to detect test-time unknown classes outside the score of trained classes. Section IV further introduces novel class discovery, including novel category discovery and its generalized setting, which requires further clustering of the new classes. Section V elaborates on continual learning. Finally, Section VI outlines future directions and Section VII concludes this paper.
Figure 3: Example applications of open-world machine learning systems. (A) Autonomous driving can leverage open-world learning to handle unknown objects and evolving environments, enabling safe navigation, adaptation to dynamic environments. (B) Medical diagnosis can encounter diverse input distributions from clients, and open-world learning enables more reliable clinical decisions and responsiveness. (C) AI chatbot continually learns from user-specific information and improves the dialogue behavior over time.
II Background and Overview
II-A Basic Notations
@@
Here, we give a holistic problem formulation and notations of open-world machine learning, comprising three tasks: unknown rejection, novel class discovery and continual learning. Let $\mathcal{D}_{\text{train}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}_{\text{train}}$ denote the training dataset, and $\mathcal{D}_{\text{test}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{M}\subset\mathcal{X}\times\mathcal{Y}_{\text{test}}$ denote the test dataset. By default, the training data only contains labeled data from old and known classes, i.e., $\mathcal{C}_{\text{old}}=\mathcal{Y}_{\text{train}}$. We represent a deep neural network (DNN) based model $f$ with two components: a feature extractor $h$ and a unified classifier $g$, i.e., $f=g\circ h$. In the canonical closed-world setting, the test dataset shares categories with training data with no additional classes, i.e., $\mathcal{Y}_{\text{test}}=\mathcal{Y}_{\text{train}}$. By contrast, in the open world, samples from new classes emerge, i.e., $\mathcal{Y}_{\text{train}}\neq\mathcal{Y}_{\text{test}}$, we further express it as $\mathcal{Y}_{\text{train}}\subset\mathcal{Y}_{\text{test}}$. In this paper, we not only consider how to classify the seen and known classes, but also focus on how to handle samples from new classes $\mathcal{C}_{\text{new}}=\mathcal{Y}_{\text{test}}\setminus\mathcal{Y}_{\text{train}}$, which is divided into three progressive ways: unknown rejection, novel class discovery and continual learning. In unknown rejection, models only need to reject unknowns, while in novel class discovery and continual learning, models are further required to classify/cluster samples from new classes. Considering the clustering objective, in novel class discovery, $\mathcal{D}_{\text{train}}$ could also contain some unlabeled data from $\mathcal{C}_{\text{new}}$. While in continual learning, the training scheme could be divided into several stages, models are trained on streaming data, and we add a superscript $t$ to the dataset $\mathcal{D}_{\text{train}}^{t}$ and $\mathcal{D}_{\text{test}}^{t}$ to indicate each continua step. Considering the differences among the three tasks, we will also give task-specific notations for the three tasks when discussing them.
II-B Canonical Use Cases
@@
In this section, we discuss several canonical applications of open-world machine learning and illustrate them in Fig. 3.
@@
1) Autonomous Driving. Self-driving cars in real-world scenarios can encounter a nearly infinite number of unexpected and novel objects [243], and open-world learning can provide them with the adaptability needed to navigate safely and effectively. Specifically, when encountering a novel object on the road, e.g., a deer or even a hole on the road that does not appear in the training class set, the car should stop or slow down. This involves leveraging an unknown rejection algorithm that provides reliable confidence scores. If the confidence score is lower than a threshold, the system should raise an alarm so that the car takes a predefined safety action. Then, novel class discovery techniques are used to create the class identity for the unknown, new object. Finally, based on continual learning methods, the car can continue to learn from a single or few samples of the object, enabling autonomous systems to quickly adapt to unexpected new environments without forgetting the skills already mastered. In conclusion, open-world learning allows the vehicle to learn from dynamic real-world environments and adapt its decision-making process, improving its driving behavior over time.
@@
2) Medical Diagnosis. Real-time medical applications require the AI system to address many heterogeneous problems that involve multiple, complex tasks. For example, a pulmonary model trained on existing lung data would not be able to reliably predict newly emerged pulmonary diseases like COVID-19, and the medical diagnostic tools can easily make mistakes [244]. In such cases, the model should be equipped with the unknown rejection ability to output reliable and human-interpretable confidence scores along with its predictions, which helps doctors or participants to make safe decisions. For example, the input should be handed over to human doctors when the confidence of a disease diagnosis is low. When encountering more and more unknowns, the medical diagnosis system is supposed to continually learn from those unknowns or more data to improve its performance progressively [245, 246]. Since manual image grading is time-consuming, novel class discovery methods would largely accelerate the processing of labeling new data. The above adaptive capability ensures that medical diagnosis systems can learn from on-the-job experiences and apply this knowledge to similar situations in the future, which is very similar to the ways that human clinicians learn. In conclusion, OWL enables medical diagnoses to make more reliable clinical decisions and be more responsive to evolving environments.
@@
3) AI Chatbot. With the launch of large language models like ChatGPT [247], AI Chatbots have become widely used in many scenarios to help with daily tasks for users. In practice, it is critical for a Chatbot to know what it doesn’t know. Specifically, if the Chatbot does not know the answer, it is better to reject to response rather than provide incorrect information that may cause very serious mistakes in fields like clinical and legal matters [248, 249]. One way to avoid providing misleading information is to abstain from making a prediction based on unknown rejection methods. Further, if the user provides the correct answer or more data is available, the Chatbot should improve its conversational skills continually without forgetting previously learned knowledge. A typical example is the hotel guest-greeting bot [7], where a Chatbot greets known hotel guests, detects new guests, and then continually learns them by asking for their names and taking photos. As can be seen, open-world learning makes the AI Chatbot more knowledgeable and powerful over time.
II-C Overall Challenges of Open-world Learning